OpenStudy (anonymous):

MEDALS AWARDED!!! An elliptical riding path is to be built on a rectangular piece of property as shown below. The rectangular piece of property measures 8 mi by 6 mi. Find an equation for the ellipse if the path is to touch the center of the property line on all 4 sides.

OpenStudy (anonymous):



OpenStudy (anonymous):

@agent0smith help plz?

OpenStudy (agent0smith):

\[\Large \frac{ x^2 }{ a^2 } + \frac{ y^2 }{b^2 } = 1\] a will be the minor axis in this case... ie distance from the centre to the horizontal edge. b is the major axis - distance from the center to the vertical edge.

OpenStudy (agent0smith):



OpenStudy (anonymous):

a=3 and b=4?

OpenStudy (anonymous):

@agent0smith ??

OpenStudy (agent0smith):

Yeah that looks right. Now use the formula from above.